OpenStudy (anonymous):

Multiply and simplify. Assume that all expressions under the radicals represent nonnegative numbers. √3x^7 √18x

OpenStudy (anonymous):

\[\sqrt a \times \sqrt b = \sqrt {ab}\]

OpenStudy (anonymous):

I got until that \[\sqrt{54x^8}\]

OpenStudy (anonymous):

and I got stuck

OpenStudy (anonymous):

what is the square root of x^8?

OpenStudy (anonymous):

2 I'm confuse with the square root and the whole thing

OpenStudy (anonymous):

what happens when you square the number 4?

OpenStudy (anonymous):

I don't really know what that means.. Sorry I'm kind of lost with this lesson

OpenStudy (anonymous):

Recall the following identity for constants a and b:\[(x^a)^b = \underbrace{\left(\underbrace{x\times x \cdots}_{a \text{ times}}\right)\left(\underbrace{x\times x \cdots}_{a \text{ times}}\right)\cdots }_{b\text{ times}}=x^{ab}\]Now, apply this to the fact that\[\sqrt{x} = x^{\frac{1}{2}}\]

OpenStudy (anonymous):

I gosh... I feel like you guys are talking japanese

OpenStudy (anonymous):

Sorry...that got cut off. The right hand side says \[(x^a)^b = x^{ab}\]

OpenStudy (anonymous):

familiapr - what is confusing to you?

OpenStudy (anonymous):

the whole thing...I don't know what the square root is or how to get the square root of a number

OpenStudy (anonymous):

okay. familiapr, do you know what the area of a square is? say the side is of length 4

OpenStudy (anonymous):

A square root is exponentiation to 1/2. Apply the formula I gave where your b=1/2. One way to think of a square root is that it outputs a number when squared will produce the original number. For instance\[\sqrt{64}=64^{\frac{1}{2}}=8\] because \[8^2=64\]

OpenStudy (anonymous):

ok got that part

OpenStudy (anonymous):

Now, you have\[\sqrt{54 x^8}\]yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

By definition, this is the same as \[\left(54x^8\right)^{\frac{1}{2}}\]and as such will abide by all normal properties of exponents. Such as that we can split them up.\[\left(54\right)^{\frac{1}{2}}\left(x^8\right)^{\frac{1}{2}}\]Of course, this is the same as also saying\[\sqrt{54}\sqrt{x^8}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

How can we simplify the \[\sqrt{x^8}\]portion?

OpenStudy (anonymous):

No idea :) sorry

OpenStudy (anonymous):

or \[\left(x^8\right)^{\frac{1}{2}}\]

OpenStudy (anonymous):

You have to use the formula I gave you.

OpenStudy (anonymous):

ok let see

OpenStudy (anonymous):

x^4

OpenStudy (anonymous):

I really don't have any idea what I'm doing

OpenStudy (anonymous):

Exactly. You can check by squaring it.\[(x^4)^2 = (x\cdot x \cdot x \cdot x)(x\cdot x \cdot x \cdot x)\]\[(x^4)^2=x^8\ \checkmark\]

OpenStudy (anonymous):

Oh gosh I'm so confuse

OpenStudy (anonymous):

Do you agree that if \[\sqrt a = b\]then\[b^2=a\]?

OpenStudy (anonymous):

LOL YOu have no idea how my brain is going around and around

OpenStudy (anonymous):

yakey, before you go any further, read what familiapr has posted earlier. he does not understand the concept of squares or square roots. you may be better served by explaining that to him.

OpenStudy (anonymous):

Can you break down the problem for me? and explain the process

OpenStudy (anonymous):

I explained what a square root is...that's just what I explained before in math symbols. The exact simplification process is as follows. Tell me what step(s) trip(s) you up.\[\sqrt{3x^7}\sqrt{18x}\]\[\sqrt{(3x^7)(18x)}\]\[\sqrt{(3\cdot 18)(x\cdot x^7)}\]\[\sqrt{54x^8}\]\[\sqrt{54}\sqrt{x^8}\]\[\left(54\right)^{1/2}\left(x^8\right)^{1/2}\]\[\left(54\right)^{1/2}x^4\]\[x^4\sqrt{54}\]

OpenStudy (anonymous):

You could even go further as follows.\[x^4\sqrt{9\times 6}\]\[x^4 \sqrt 9 \sqrt 6\]\[3x^4 \sqrt 6\]

OpenStudy (anonymous):

ok can explain me why x^8 doesn't stay like that?

OpenStudy (anonymous):

\[\sqrt{x^8}\]\[(x^8)^{\frac 12}\]\[x^{8\cdot\frac{1}{2}}\]\[x^4\]

OpenStudy (anonymous):

I get that , what I'm asking is when you split that x^8 into X^7 and x

OpenStudy (anonymous):

I didn't split them...I combined them? They were given separate in the problem, so to simplify you combine them? I was starting from the original problem.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Can you help me with this one? I try and I don't know if is correct I get stuck \[\sqrt[4]{48x^4y^6}\]